A stone is tossed upward at the moment a ball is dropped from a height of 25 meters. The stone’s initial velocity is 25 m/s. At what height will the two meet?

 

 



1
Expert's answer
2021-09-29T09:56:04-0400

The equations of motion of the objects are given by


"y_{\rm stone}=v_0t-gt^2/2=25t-4.9t^2""y_{\rm ball}=h-gt^2/2=25-4.9t^2"

At the instant when objects meet

"y_{\rm stone}=y_{\rm ball}"

Hence

"25t-4.9t^2=25-4.9t^2"

"t=1\:\rm s"

"y_{\rm stone}=y_{\rm ball}=25-4.9*1^2=20.1\:\rm m"